how do you work out
x^2-4y^2-2x-24y-39=0

Answer by nerdybill(7384)

x^2-4y^2-2x-24y-39=0
reorder terms:
x^2-2x-4y^2-24y = 39
group terms:
(x^2-2x)-(4y^2+24y) = 39
(x^2-2x)-4(y^2+6y) = 39
complete squares:
(x^2-2x+1)-4(y^2+6y+9) = 39+1-36
(x-1)^2-4(y+3)^2 = 4
(x-1)^2 /4 - (y+3)^2 = 1
.
This is a horizontal hyperbola
(h,k) is (1,-3)
a is 2
b is 1